Please help me, I need to get through this test to finish school.
1. Multiply and Simplify
4y^2/5y^2-40y+80 * 5y-20/2y =
(Type and Exponential notation with positive exponents)
2. Multiply
(8x+6)(5x^2+5x+8)
3. Evaluate the Polynomial for x=6
2x^2-5x+7=
4. Solve.
(x+19)(x-2)(x+10)>0
Select the correct choice below and if necessary, fill in the answer box to complete your choice.
A. The solution set is {|x| }
B. The solution is all real numbers
C. There is no solution.


Answer by stanbon(75887)

1. Multiply and Simplify
4y^2/(5y^2-40y+80) * (5y-20)/2y
-------------
= (4y^2)/[5(y-4)^2] * (5(y-4))/(2y)
----
Cancel where you can to get:
= (2y)/(y-4)
======================================
2. Multiply
(8x+6)(5x^2+5x+8)
----
= (8x)(5x^2+5x+8) + 6(5x^2+5x+8)
----
= 40x^3+40x^2+64x + 30x^2+30x+48
----
= 40x^3+70x^2+94x+48
======================================

3. Evaluate the Polynomial for x=6
2x^2-5x+7= 2(36)-5(6)+7 = 72-30+7 = 49
===========================================
4. Solve.
(x+19)(x-2)(x+10)>0
Draw a number line.
Plot x = -19 ; x = -10 ; x = 2
Test a value in each of the 4 intervals to see where the solutions are.
(x+19)(x-2)(x+10)>0
--
Test x = -25 to get: -*-*- > 0; false
Test x = -15 to get: +*-*- > 0; true so solutions in(-19,-10)
Test x = 0 to get:: +*-*+ > 0; false
Test x = 5 to get:: +*+*+ > 0; true so solutions in (2,+oo)
---
Solution: {x | -19< x < -10 or 2< x <+oo}
==============================================
Cheers,
Stan H.
==============

=============
Select the correct choice below and if necessary, fill in the answer box to complete your choice.
A. The solution set is {|x| }
B. The solution is all real numbers
C. There is no solution.